I am drawing blanks on remember how to factor and check answers to make sure they are correct. 2 examples are 4x^2=12x and 13x^2-66x+5=0..could someone please help me remember how to do this!!!
Answer by checkley77(12844)

4x^2=12x
4x^2-12x=0 factor out the common terms 4x
4x(x-3)=0
4x=0
x=0 ans.
x-3=0
x=3 ans.
Proofs:
4*0^2=12*0
0=0
4*3^2=12*3
4*9=36
36=36
--------------------
13x^2-66x+5=0
(13x-1)(x-5)=0
13x-1=0
13x=1
x=1/13 ans.
x-5=0
x=5 ans.
Proofs:
13(1/13)^2-66*1/13=5=0
13*1/13*13-66/13+5=0
1/13-66/13+5=0
-65/13+5=0
-5+5=0
0=0